What cycle time would match capacity and demand if demand is 172 units a day, there are 1 shift(s) of 457 minutes each, and workers are given two 30 minute breaks during each shift, one of which is for lunch or dinner? (Round your answer to 2 decimal places.)

Cycle time             minutes per cycle
Answer #1

Net time available for production = 457-2*30 = 397

Cycle time = Net time available for production/Daily demand = 397/172 = 2.30814 = 2.31 minutes per cycle (Rounded to 2 decimal places)